### v3

# Are Self-Attentions Effective for Time Series Forecasting?

Abstract
Time series forecasting is crucial for applications across multiple domains and various scenarios. Although Transformers have dramatically advanced the landscape of forecasting, their effectiveness remains debated. Recent findings have indicated that simpler linear models might outperform complex Transformer-based approaches, highlighting the potential for more streamlined architectures. In this paper, we shift the focus from evaluating the overall Transformer architecture to specifically examining the effectiveness of self-attention for time series forecasting. To this end, we introduce a new architecture, Cross-Attention-only Time Series transformer (CATS), that rethinks the traditional transformer framework by eliminating self-attention and leveraging cross-attention mechanisms instead. By establishing future horizon-dependent parameters as queries and enhanced parameter sharing, our model not only improves long-term forecasting accuracy but also reduces the number of parameters and memory usage. Extensive experiment across various datasets demonstrates that our model achieves superior performance with the lowest mean squared error and uses fewer parameters compared to existing models. The implementation of our model is available at: https://github.com/dongbeank/CATS.
11footnotetext: Corresponding authors
1 Introduction
Time series forecasting plays a critical role within the machine learning society, given its applications ranging from financial forecasting to medical diagnostics. To improve the accuracy of predictions, researchers have extensively explored and developed various models. These range from traditional statistical methods to modern deep learning techniques. Most notably, Transformer [19] has brought about a paradigm shift in time series forecasting, resulting in numerous high-performance models, such as Informer [29], Autoformer [23], Pyraformer [11], FEDformer [31], and Crossformer [28]. This line of work establishes new benchmarks for high performance in time series forecasting.
However, Zeng et al. [26] have raised questions about the effectiveness of Transformer-based time series forecasting models especially for long term time series forecasting. Specifically, their experiments demonstrated that simple linear models could outperform these Transformer-based approaches, thereby opening new avenues for research into simpler architectural frameworks. Indeed, the following studies [12, 6] have further validated that these linear models can be enhanced by incorporating additional features.
Despite these developments, the effectiveness of each components in Transformer architecture in time series forecasting remains a subject of debate. Nie et al. [14] introduced an encoder-only Transformer that utilizes patching rather than point-wise input tokens, which exhibited improved performance compared to linear models. Zeng et al. [26] also highlighted potential shortcomings in simpler linear networks, such as their inability to capture temporal dynamics at change points [18] compared to Transformers. Consequently, while a streamlined architecture may be beneficial, it is imperative to critically evaluate which elements of the Transformer are necessary and which are not for time series modeling.
In light of these considerations, our study shifts focus from the overall architecture of the Transformer to a more specific question: Are self-attentions effective for time series forecasting? While this question is also noted in [26], their analysis was limited to substituting attention layers with linear layers, leaving substantial room for potential model design when focusing on Transformers. Furthermore, the issue of temporal information loss (i.e., permutation-invariant and anti-order characteristics of self-attention) is predominantly caused by the use of self-attention rather than the Transformer architecture itself. Therefore, we aim to resolve the issues of self-attention and therefore propose a new forecasting architecture that achieves higher performance with a more efficient structure.
Figure 1: Experimental results illustrating the mean squared error (MSE) and the number of parameters with varying input sequence lengths on ETTm1. Each bubble represents a different model, with the bubble size indicating the number of parameters in millions—larger bubbles denote models with more parameters. Our model consistently shows the lowest MSE (i.e., best performance) with fewer parameters even for longer input sequences. The detailed results can be found in Table 3.
In this paper, we introduce a novel forecasting architecture named Cross-Attention-only Time Series transformer (CATS) that simplifies the original Transformer architecture by eliminating all self-attentions and focusing on the potential of cross-attentions. Specifically, our model establishes future horizon-dependent parameters as queries and treats past time series data as key and value pairs. This allows us to enhance parameter sharing and improve long-term forecasting performance. As shown in Figure 1, our model shows the lowest mean squared error (i.e., better forecasting performance) even for longer input sequences and with fewer parameters than existing models. Moreover, we demonstrate that this simplified architecture can provide a clearer understanding of how future predictions are derived from past data with individual attention maps for the specific forecasting horizon. Finally, through extensive experiments, we show that our proposed model not only achieves state-of-the-art performance but also requires fewer parameters and less memory consumption compared to previous Transformer models across various time series datasets.
2 Related Work
Time Series Transformers
Transformer models [19] have shown effective in various domains [5, 4, 15], with a novel encoder-decoder structure with self-attention, masked self-attention, and cross-attention. The self-attention mechanism is a key component for extracting semantic correlations between paired elements, even with identical input elements; however, autoregressive inference with self-attention requires quadratic time and memory complexity. Therefore, Informer [29] proposed directly predicting multi-steps, and a line of work, such as Autoformer [23], FEDformer [31], and Pyraformer [11], investigated the complexity issue in time series transformers. Simultaneously, unique properties of time series, such as stationarity [12], decomposition [23], frequency features [31], or cross-dimensional properties [28] were employed to modify the attention layer for forecasting tasks. Recently, researchers have investigated the essential architecture in Transformers to capture long-term dependencies. PatchTST [14] became a de-facto standard Transformer model by patching the time series input in a channel-independence manner, which is widely used in following Transformer-based forecasting models [13, 7]. On the other hand, Das et al. [3] emphasized the importance of decoder-only forecasting models, while they focused on zero-shot using pre-trained language models. However, none of them have investigated the importance of cross-attention for time series forecasting.
Temporal Information Encoding
Fixed temporal order in time series is the distinct property of time series, in contrast to the language domain where semantic information does not heavily depend on the word ordering [4]. Thus, some researchers have used learnable positional encoding in Transformers to embed time-dependent properties [9, 23]. However, Zeng et al. [26] first argued that self-attention is not suitable for time series due to its permutation invariant and anti-order properties. While they focus on building complex representations, they are inefficient in maintaining the original context of historical and future values. They rather proposed linear models without any embedding layer and demonstrated that it can achieve better performance than Transformer models, particularly showing robust performance to long input sequences. Recent linear time series models outperformed previous Transformer models with simple architectures by focusing on pre-processing and frequency-based properties [10, 2, 21]. On the other hand, Woo et al. [22] investigated the new line of works of time-index models, which try to model the underlying dynamics with given time stamps. These related works imply that preserving the order of time series sequences plays a crucial role in time series forecasting.
3 Revisiting Self-Attention in Time Series Forecasting
Motivation of Self-Attention Removal
Following the concerns about the effectiveness of self-attention on temporal information preservation [26], we conduct an experiment using PatchTST [14]. We consider three variations of the PatchTST model: the original PatchTST with overlapping patches with length 16 and stride 8 (Fig. 2a); a modified PatchTST with non-overlapping patches with length 24 (Fig. 2b); and a version where self-attention is replaced by a linear embedding layer, using non-overlapping patches with length 24 (Fig. 2c). This setup allows us to isolate the effects of self-attention on temporal information preservation, while controlling for the impact of patch overlap.
(a) Original PatchTST
(b) PatchTST w/ non-overlapping
(c) PatchTST w/o self-attn
Figure 2: Absolute values of weights in the final linear layer for different PatchTST variations. The distinct patterns reveal how each model captures temporal information.
Fig. 2 illustrates the absolute values of the weights in the final linear layer for these model variations. Compared to the original PatchTST (Fig. 2a), both non-overlapping versions (Fig. 2b and Fig. 2c) show more vivid patterns. The version with linear embedding (Fig. 2c) demonstrates the clearest capture of temporal information, suggesting that the self-attention mechanism itself may not be necessary for capturing temporal information.
In Table 3, we summarize the forecasting performance of the original PatchTST (Fig. 2a) and PatchTST without self-attention (Fig. 2c). PatchTST without self-attention consistently improves or maintains performance across all forecasting horizons. Specifically, the original version with self-attention shows lower performance for longer forecast horizons. This result suggests that self-attention may not only be unnecessary for effective time series forecasting but could even hinder
[…]
tableEffect of self-attention in PatchTST on forecasting performance (MSE) on ETTm1. Horizon original w/o self-attn 96 0.290 0.290 192 0.332 0.328 336 0.366 0.359 720 0.416 0.414
performance in certain cases. Therefore, better performance of w/o self-attention challenges the conventional belief regarding the importance of self-attention mechanisms in Transformer-based models for time series forecasting tasks.
Our findings offer new insights into the role of self-attention in time series forecasting. As shown in Fig. 2 and Table 3, replacing self-attention with a linear layer not only captures clear temporal patterns but also results in significant performance improvements, particularly for longer forecast horizons. These results highlight potential areas for enhancing the handling of temporal information, beyond addressing the well-known concerns regarding computational complexity.
Rethinking Transformer Design
(a) Transformer
(b) Encoder
(c) Linear model
(d) Ours
Figure 3: Illustration of existing time series forecasting architectures and the proposed architecture.
Given the challenges associated with self-attention in time series forecasting, we propose a fundamental rethinking of the Transformer architecture for this task. Fig. 3 illustrates the differences between existing architectures and our proposed approach. Traditional Transformer architectures (Fig. 3a) and encoder-only models (Fig. 3b) rely heavily on self-attention mechanisms, which may lead to temporal information loss. In contrast, Zeng et al. [26] proposed a simplified linear model, DLinear (Fig. 3c), which removes all Transformer-based components. While this approach reduces computational load and potentially avoids some temporal information loss, it may struggle to capture complex temporal dependencies.
To address these challenges while preserving the advantages of Transformer architectures, we propose the Cross-Attention-only Time Series transformer (CATS), depicted in Fig. 3d. Our approach removes all self-attention layers and focuses solely on cross-attention, aiming to better capture temporal dependencies while maintaining the structural advantages of the transformer architecture. In the following section, we will introduce our CATS model in detail, explaining our key innovations including a novel use of cross-attention, efficient parameter sharing, and adaptive masking techniques.
4 Proposed Methodology
4.1 Problem Definition and Notations
A multivariate time series forecasting task aims to predict future values $\bm{\tilde{X}}=\{{\mathbf{x}}_{L+1},\ldots,{\mathbf{x}}_{L+T}\}\in\mathbb{R}^{M\times T}$ with the prediction $\bm{\hat{X}}=\{\hat{{\mathbf{x}}}_{L+1},\ldots,\hat{{\mathbf{x}}}_{L+T}\}\in\mathbb{R}^{M\times T}$ based on past datasets $\bm{X}=\{{\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{L}\}\in\mathbb{R}^{M\times L}$. Here, $T$ represents the forecasting horizon, $L$ denotes the input sequence length, and $M$ represents the dimension of time series data.
In traditional time series transformers, we feed the historical multivariate time series $\bm{X}$ to embedding layers, resulting in the historical embedding $\bm{H}\in\mathbb{R}^{D\times L}$. Here, $D$ is the embedding size. Note that, in channel-independence manners, the multivariate input is considered to separate univariate time series ${\mathbf{x}}\in\mathbb{R}^{1\times L}$. With patching [14], univariate time series ${\mathbf{x}}$ transforms into patches ${\mathbf{p}}=\texttt{Patch}({\mathbf{x}})\in\mathbb{R}^{P\times N_{L}}$ where $P$ is the size of each patch and $N_{L}$ is the number of input patches. Similar to non-patching situations, patches are fed to embedding layers $\bm{P}=\texttt{Embedding}({\mathbf{p}})\in\mathbb{R}^{D\times N_{L}}$.
Figure 4: Illustration on the proposed model architecture. Our model removes all self-attentions from the original Transformer structure and focuses on cross-attentions. To fully utilize the cross-attention, we conceptualize the future horizon as queries and use the input time series (i.e., past time series) as keys and values (Fig. A). This simplified structure enables us to enhance the parameter sharing across forecasting horizons (Fig. B) and make use of query-adaptive masking (Fig. C) for performance.
4.2 Model Structure
Our proposed architecture consists of three key components: (A) Cross-Attention with Future as Query, (B) Parameter Sharing across Horizons, and (C) Query-Adaptive Masking. Fig. 4 illustrates how our model modifies the traditional transformer structure.
By focusing solely on cross-attention, our approach allows us to maintain the periodic properties of time series, which self-attention, with its permutation-invariant and anti-order characteristics, struggles to capture. Furthermore, we leverage advanced architectural designs of time series transformers, such as patching [14], which linear models cannot utilize. The following subsections provide detailed descriptions of each component of our CATS model, explaining how these elements work together to address the challenges of time series forecasting.
Cross-Attention via Future as Query
Similar to self-attention, the cross-attention mechanism employs three elements: key, query, and value. The distinctive feature of cross-attention is that the query originates from a different source than the key or value. Generally, the query component aims to identify the most relevant information among the keys and uses it to extract crucial data from the values [1, 27]. In the realm of time series forecasting, where predictions are often made for a specific target horizon—such as forecasting 10 steps ahead. Therefore, within this concept of forecasting, we argue that each future horizon should be regarded as a question, i.e., an independent query.
To implement this, we establish horizon-dependent parameters as learnable queries. As shown in Fig. 4, we begin by creating parameters for the specified forecasting horizon. For each of these virtualized parameters, we assign a fixed number of parameters to represent the corresponding horizon as learnable queries ${\mathbf{q}}$. For example, ${\mathbf{q}}_{i}$ is a horizon-dependent query at $L+i$. When patching is applied, these queries are then processed independently; each learnable query ${\mathbf{q}}\in\mathbb{R}^{P}$ is first fed into the embedding layer, and then fed into the multi-head attention with the embedded input time series patches as the key and value.
Based on these new query parameters, we can utilize a cross-attention-only structure in the decoder, resulting in an advantage in efficiency. In Table 1, we summarize the time complexity of recent transformer models and ours. The results indicate that our method only requires the time complexity of $\mathcal{O}({LT}/{P^{2}})$, where most of the transformer-based models require $\mathcal{O}({L}^{2})$ except FEDformer and Pyraformer. However, since these two models have an encoder-decoder and a relatively huge amount of parameters, they require 10x and 4x computational times than ours, respectively.
Table 1: Time complexity of transformer-based models to calculate attention outputs. Time refers to the inference time obtained by averaging 10 runs under $L=96$ and $T=720$ on Electricity.
Method Encoder Decoder Time Method Encoder Decoder Time
Transformer [13] $\mathcal{O}(L^{2})$ $\mathcal{O}(T(T+L))$ 10.4ms Informer [29] $\mathcal{O}(L\log L)$ $\mathcal{O}(T(T+\log L))$ 13.5ms
Autoformer [23] $\mathcal{O}(L\log L)$ $\mathcal{O}\left(\left(L/2+H\right)\log\left(L/2+T\right)\right)$ 24.1ms Pyraformer [11] $\mathcal{O}(L)$ $\mathcal{O}(T(T+L))$ 11.2ms
FEDformer [31] $\mathcal{O}(L)$ $\mathcal{O}\left(L/2+H\right)$ 69.3ms Crossformer [28] $\mathcal{O}\left({M}L^{2}/{P^{2}}\right)$ $\mathcal{O}\left({M}T(T+L)/{P^{2}}\right)$ 30.6ms
PatchTST [14] $\mathcal{O}\left({L^{2}}/{P^{2}}\right)$ 7.6ms CATS (Ours) - $\mathcal{O}\left({LT}/{P^{2}}\right)$ 7.0ms
Parameter Sharing across Horizons
One of the strongest benefits of cross-attention via future horizon as a query ${\mathbf{q}}$ is that each cross-attention is only calculated on the values from a single forecasting horizon and the input time series. Mathematically, for a prediction of future value $\hat{{\mathbf{x}}}_{L+i}$ can be expressed as a function solely dependent on the past samples $\bm{X}=[{\mathbf{x}}_{1},...,{\mathbf{x}}_{L}$] and ${\mathbf{q}}_{i}$, independent of ${\mathbf{q}}_{j}$ for all $i\neq j$ or $i$ and $j$ are not in the same patch.
This independent forecasting mechanism offers a notable advantage; a higher level of parameter sharing. As demonstrated in [14], significant reductions in the required number of parameters can be achieved in time series forecasting through parameter sharing between inputs or patches, enhancing computational efficiency. Regarding this, we propose parameter sharing across all possible layers — the embedding layer, multi-head attention, and projection layer — for every horizon-dependent query ${\mathbf{q}}$. In other words, all horizon queries ${\mathbf{q}}_{1},\dots,{\mathbf{q}}_{T}$ or ${\mathbf{q}}_{1},\dots,{\mathbf{q}}_{{N_{T}}}$ share the same embedding layer used for the input time series ${\mathbf{x}}_{1},\dots,{\mathbf{x}}_{L}$ or patches ${\mathbf{p}}_{1},\dots,{\mathbf{p}}_{N_{L}}$ before proceeding to the cross-attention layer, respectively. Furthermore, to maximize the parameter sharing, we also propose cross-dimension sharing that uses the same query parameters for all dimensions.
For the multi-head attention and projection layers, we apply the same algorithm across horizons. Notably, unlike the approach in PatchTST [14], we also share the projection layer for each prediction. Specifically, PatchTST, being an encoder-only model, employs a fully connected layer as the projection layer for the encoder outputs $\bm{P}\in\mathbb{R}^{D\times N_{L}}$, resulting in $(D\times N_{L})\times T$ parameters. In contrast, our model first processes raw queries ${\mathbf{q}}=[{\mathbf{q}}_{1},\ldots,{\mathbf{q}}_{N_{T}}]\in\mathbb{R}^{P\times N_{T}}$. These queries are then embedded through the cross-attention mechanism, resulting in $\bm{Q}=[\bm{q}_{1},\ldots,\bm{q}_{N_{T}}]\in\mathbb{R}^{D\times N_{T}}$. The final projection uses shared parameters $W\in\mathbb{R}^{P\times D}$, producing an output $W\bm{Q}\in\mathbb{R}^{P\times N_{T}}$. Thus, our number of parameters for this projection becomes $P\times D$, which is not proportionally increasing to $T$. This approach significantly reduces time complexity during both the training and inference phases.
[…]
tableEffect of parameter sharing across horizons on the number of parameters for different forecasting horizons on ETTh1. Horizon w/ sharing w/o sharing 96 355,320 404,672 192 355,416 552,320 336 355,560 958,112 720 355,944 3,121,568
In Table 4.2, we outline the impact of parameter sharing across different forecasting horizons. In contrast to the model without parameter sharing, which exhibits a rapid increase in parameters as the forecasting horizon extends, our model, which shares all layers including the projection layer, maintains a nearly consistent number of parameters.
Additionally, all operations, including embedding and multi-head attention, are performed independently for each learnable query. This implies that the forecast for a specific horizon does not depend on other horizons. Such an approach allows us to generate distinct attention maps for each forecasting horizon, providing a clear understanding of how each prediction is derived. Please refer to Section 5.5.
Query-Adaptive Masking
Parameter sharing across horizons enhances the efficiency of our proposed architecture and simplifies the model. However, we observed that a high degree of parameter sharing could lead to overfitting to the keys and values (i.e., past time series data), rather than the queries (i.e., forecasting horizon). Specifically, the model may converge to generate similar or identical predictions, $\hat{{\mathbf{x}}}_{L+i}$ and $\hat{{\mathbf{x}}}_{L+j}$, despite receiving different horizon queries, ${\mathbf{q}}_{i}$ and ${\mathbf{q}}_{j}$ (i.e., the target horizons differ).
Therefore, to ensure the model focuses on each horizon-dependent query ${\mathbf{q}}$, we introduce a new technique that masks the attention outputs. As illustrated in the right-bottom figure of Fig. 4, for each horizon, we apply a mask to the direct connection from Multi-Head Attention to LayerNorm with a probability $p$. This mask prevents access to the input time series information, resulting in only the query to influence future value predictions. This selective disconnection, rather than the application of dropout in the residual connections, helps the layers to concentrate more effectively on the forecasting queries. We note that this approach can be related to stochastic depth in residual networks [8]. The stochastic depth technique has proven effective across various tasks, such as vision tasks [17, 25]. To the best of our knowledge, this is the first application of stochastic depth in Transformers for time series forecasting. A detailed analysis of query-adaptive masking can be found in Appendix.
In summary, the framework described in this section, including cross-attention via future as query, parameter sharing across horizons, and query-adaptive masking, is named the Cross-Attention-only Time Series transformer (CATS).
5 Experiments
In this section, we provide extensive experiments to provide the benefits of our proposed framework, CATS, including forecasting performance and computational efficiency. To this end, we use 7 different real-world datasets and 9 baseline models. For datasets, we use Electricity, ETT (ETTh1, ETTh2, ETTm1, and ETTm2), Weather, Traffic, and M4. These datasets are provided in [23] and [24] for time series forecasting benchmark, detailed in Appendix.
For baselines, we utilize a wide range of various baselines, including the state-of-the-art long-term time series forecasting model TimeMixer [21], PatchTST [14], Timesnet [24], Crossformer [28], MICN [20], FiLM [30], DLinear [26], Autoformer [23], and Informer [29]. For both long-term and short-term time series forecasting results, we report performance of our model alongside the results of other models as presented in TimeMixer [21], ensuring a consistent comparison across all baselines. We used 4 NVIDIA RTX 4090 24GB GPUs with 2 Intel(R) Xeon(R) Gold 5218R CPUs @ 2.10GHz for all experiments.
5.1 Long-Term Time Series Forecasting Results
To ease comparison, we follow the settings of [21] for long-term forecasting, using various forecast horizons with a fixed 96 input sequence length. Detailed settings are provided in the Appendix. Table 9 summarizes the forecasting performance across all datasets and baselines. Our proposed model, CATS, demonstrates superior performance in multivariate long-term forecasting tasks across multiple datasets. CATS consistently achieves the lowest Mean Squared Error (MSE) and Mean Absolute Error (MAE) on the Traffic dataset for all forecast horizons, outperforming all other models. For the Weather, Electricity, and ETT datasets, CATS shows competitive performance, achieving the best results on most forecast horizons. This indicates that CATS effectively captures underlying patterns in diverse time series data, highlighting its capability to handle complex temporal dependencies. Additional experiments with longer input sequence lengths of 512 are provided in the Appendix.
Table 2: Multivariate long-term forecasting results with recent forecasting models and ours for unified hyperparameter settings. The best results are in bold and the second best are underlined. Full results are provided in Appendix.
Models CATS TimeMixer PatchTST Timesnet Crossformer MICN FiLM DLinear Autoformer Informer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Weather 96 0.161 0.207 0.163 0.209 0.186 0.227 0.172 0.220 0.195 0.271 0.198 0.261 0.195 0.236 0.195 0.252 0.266 0.336 0.300 0.384
192 0.208 0.250 0.208 0.250 0.234 0.265 0.219 0.261 0.209 0.277 0.239 0.299 0.239 0.271 0.237 0.295 0.307 0.367 0.598 0.544
336 0.264 0.290 0.251 0.287 0.284 0.301 0.246 0.337 0.273 0.332 0.285 0.336 0.289 0.306 0.282 0.331 0.359 0.395 0.578 0.523
720 0.342 0.341 0.339 0.341 0.356 0.349 0.365 0.359 0.379 0.401 0.351 0.388 0.361 0.351 0.345 0.382 0.419 0.428 1.059 0.741
Electricity 96 0.149 0.237 0.153 0.247 0.190 0.296 0.168 0.272 0.219 0.314 0.180 0.293 0.198 0.274 0.210 0.302 0.201 0.317 0.274 0.368
192 0.163 0.250 0.166 0.256 0.199 0.304 0.184 0.322 0.231 0.322 0.189 0.302 0.198 0.278 0.210 0.305 0.222 0.334 0.296 0.386
336 0.180 0.268 0.185 0.277 0.217 0.319 0.198 0.300 0.246 0.337 0.198 0.312 0.217 0.300 0.223 0.319 0.231 0.443 0.300 0.394
720 0.219 0.302 0.225 0.310 0.258 0.352 0.220 0.320 0.280 0.363 0.217 0.330 0.278 0.356 0.258 0.350 0.254 0.361 0.373 0.439
Traffic 96 0.421 0.270 0.462 0.285 0.526 0.347 0.593 0.321 0.644 0.429 0.577 0.350 0.647 0.384 0.650 0.396 0.613 0.388 0.719 0.391
192 0.436 0.275 0.473 0.296 0.522 0.332 0.617 0.336 0.665 0.431 0.589 0.356 0.600 0.361 0.598 0.370 0.616 0.382 0.696 0.379
336 0.453 0.284 0.498 0.296 0.517 0.334 0.629 0.336 0.674 0.420 0.594 0.358 0.610 0.367 0.605 0.373 0.622 0.337 0.777 0.420
720 0.484 0.303 0.506 0.313 0.552 0.352 0.640 0.350 0.683 0.424 0.613 0.361 0.691 0.425 0.645 0.394 0.660 0.408 0.864 0.472
ETT (Avg) 96 0.289 0.339 0.290 0.339 0.326 0.362 0.312 0.355 0.465 0.456 0.340 0.388 0.324 0.358 0.319 0.368 0.389 0.415 1.414 0.816
192 0.348 0.374 0.350 0.373 0.388 0.397 0.365 0.385 0.553 0.518 0.408 0.431 0.384 0.393 0.399 0.418 0.448 0.443 1.985 0.989
336 0.376 0.395 0.390 0.404 0.426 0.423 0.455 0.421 0.686 0.584 0.479 0.476 0.428 0.423 0.469 0.463 0.491 0.473 2.101 1.101
720 0.434 0.433 0.439 0.438 0.464 0.455 0.467 0.455 1.038 0.754 0.597 0.541 0.481 0.459 0.596 0.537 0.533 0.504 2.343 1.163
5.2 Efficient and Robust Forecasting for Long Input Sequences
Zeng et al. [26] observed that many models experience a decline in performance when using long input sequences for time series forecasting. To address this, some approaches have been developed to capture long-term dependencies. For instance, TimeMixer [21] employs linear models with mixed scale, and PatchTST [14] utilizes an encoder network to encode long-term information. However, these models still have computational issues, particularly in terms of escalating memory and parameter requirements. Thus, in this subsection, we provide a comparison between previous models and ours in terms of efficient and robust forecasting for long input sequences.
First of all, to provide a fair comparison, we summarize the number of parameters, GPU memory consumption, and forecasting performance of comparison models with varying input lengths. As summarized in Table 3, existing complex models, such as PatchTST and TimeMixer, suffer from increased parameters and computational burdens when performing forecasting with long input lengths. Although DLinear uses fewer parameters and less GPU memory, its performance is limited due to its linear structure in capturing non-linearity patterns. Considering both performance and efficiency, the proposed model demonstrates robust performance improvement even with longer input lengths. In Appendix, we provide additional experimental results supporting these findings.
Table 3: Comparison of models with the number of parameters, GPU memory consumption, and MSE across different input sequence lengths on ETTm1. Full results with more diverse input lengths are provided in Appendix.
Parameters GPU Memory MSE
Input Length 336 720 1440 2880 336 720 1440 2880 336 720 1440 2880
PatchTST 4.3M 8.7M (2.0x) 17.0M (4.0x) 33.6M (7.9x) 3.5GB 7.4GB (2.1x) 22.0GB (6.3x) 58.6GB (16.9x) 0.418 0.418 0.420 0.412
TimeMixer 1.1M 4.1M (3.6x) 14.2M (12.6x) 52.9M (46.8x) 2.9GB 3.9GB (1.3x) 5.9GB (2.0x) 10.3GB (3.6x) 0.428 0.425 0.414 0.472
DLinear 0.5M 1.0M (2.1x) 2.1M (4.2x) 4.2M (8.5x) 1.1GB 1.1GB (1.0x) 1.2GB (1.0x) 1.2GB (1.1x) 0.426 0.422 0.401 0.408
CATS 0.4M 0.4M (1.0x) 0.4M (1.0x) 0.4M (1.1x) 1.9GB 2.1GB (1.1x) 2.7GB (1.4x) 3.8GB (2.0x) 0.407 0.402 0.399 0.395
(a) Model Performance
(b) Parameter Efficiency
(c) Memory Efficiency
(d) Running Time Efficiency
Figure 5: Efficiency and performance analysis for time series forecasting models. We summarize the forecasting performance, number of parameters, GPU memory consumption, and running time with varying forecasting horizon lengths on Traffic. The running time is averaged from 300 iterations.
Furthermore, we conduct a deeper comparison between transformer-based models. Especially, TimeMixer [21] argues that their model outperforms PatchTST [14] in the setting of long input sequences. Regarding this setting, we also conduct an experiment on $L=512$. We summarize the results in Fig. 5. Among these transformer-based models, our model achieves the lowest MSE for most forecasting horizons. Moreover, our model requires even a lower number of parameters, GPU memory, and running time. Especially, for parameter efficiency, CATS shows significant differences even on a log scale due to its efficient parameter-sharing. Fig. 5c highlights GPU memory usage across different forecasting horizons. While PatchTST and TimeMixer consume significantly more memory, CATS maintains a low and stable memory consumption, demonstrating superior memory efficiency. In Fig. 5d CATS also consistently achieves lower running times compared to PatchTST and TimeMixer.
Additionally, we also compare the same factors when we use a longer input length $L=2880$. As more input length is used, the forecasting performance of our model outperforms all other models. Most importantly, while the computational complexity increases as input length increases, our model achieves a faster running time, compared to other models with a 512 input sequence length. Overall, these results emphasize the efficiency and performance advantages of our model, particularly in terms of parameter count, memory usage, and running time.
5.3 Short-Term Time Series Forecasting Results
Table 4 summarizes the averaged results for the M4 dataset, comparing the performance of various models. Our CATS model consistently achieved the best results across all metrics. Notably, CATS reduced MASE by 19.94% compared to PatchTST, a self-attention-only model that suffers from temporal information loss as mentioned in Section 3. CATS, with its cross-attention-only structure, effectively mitigated this issue and captured temporal dependencies more efficiently than previous models. Although TimeMixer, a state-of-the-art linear model, performed well, CATS surpassed it across all metrics. This demonstrates that CATS excels at capturing short-term temporal dependencies, providing superior performance in short-term forecasting tasks.
Table 4: Averaged univariate short-term forecasting results in the M4 dataset. The best results are in bold and the second best are underlined. Full results are presented in Appendix.
Models CATS TimeMixer Timesnet PatchTST MICN FiLM DLinear Autoformer Informer
Average SMAPE 11.701 11.723 11.829 13.152 19.638 14.863 13.639 12.909 14.086
MASE 1.557 1.559 1.585 1.945 5.947 2.207 2.095 1.771 2.718
OWA 0.838 0.840 0.851 0.998 2.279 1.125 1.051 0.939 1.230
5.4 Replacement of Cross-attention with Self-attention
In our propose structure, we mainly use cross-attention rather than self-attention due to the forecasting-unfriendly properties of self-attention. To verify the effectiveness of cross-attention in the proposed structure, we replace the cross-attention layers with self-attention layers while maintaining other structures. In Table 5, we gradually replace the cross-attention with self-attention (SA) among a total of three cross-attention layers. To maintain the original transformer structure, we set the maximum replacement as two. As shown in this table, we confirm the effectiveness of the cross-attention mechanism compared to using self-attention layers. Specifically, the zero SA, which is our model, shows better performance than using SA for almost all cases except only one case.
Table 5: Performance comparison on models with three attention layers. We replace one or more cross-attentions (CA) with self-attentions (SA) in our model. In total, there are three cross-attentions in all settngs and ‘Zero SA’ stands for our model. The best results are in bold.
Dataset Electricity ETTm1
Case Zero SA One SA Two SA Zero SA One SA Two SA
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
96 0.126 0.218 0.128 0.220 0.133 0.225 0.283 0.340 0.284 0.338 0.284 0.340
192 0.144 0.235 0.150 0.238 0.153 0.245 0.319 0.363 0.331 0.373 0.324 0.368
336 0.159 0.252 0.167 0.257 0.169 0.263 0.351 0.385 0.376 0.401 0.369 0.400
720 0.194 0.283 0.205 0.293 0.210 0.300 0.400 0.414 0.429 0.437 0.442 0.445
5.5 Explaining Periodic Patterns through Cross-Attention
As noted in Section 4.2, in our proposed model, all operations including embedding and multi-head attention are performed independently for each learnable query. In other words, the forecast for a specific horizon does not depend on other horizons. This approach helps us better understand how each prediction is derived. Therefore, in this subsection, we visualize how the proposed model understands the periodic properties.
To provide an easy understanding, we here consider a simple time series forecasting task with data that consists of two independent signals as follows:
${\mathbf{x}}(t)$ =$=\{x_{(t\bmod\tau)}\}^{\infty}_{t=1},\quad x_{i}\sim\mathcal{N}(0,1)\quad(i=0,1,\ldots,\tau-1),$ (1)
${\bm{y}}(t)$ =$=\begin{cases}+k&\text{if }t\equiv 0\quad\pmod{S}\\ -k&\text{if }t\equiv\frac{1}{2}S\text{ }\pmod{S}.\end{cases}$ (2)
For prediction, we use an input sequence length $L=48$ and a forecasting horizon $T=72$ with signals ${\mathbf{x}}(t)$ and ${\bm{y}}(t)$ are defined with $\tau=24$, $S=8$, and $k=5$. The patch length is set to 4 without overlapping to elucidate the distinct periodic components with 2 attention heads.
(a) Cross-attention score map of head 1
(b) Cross-attention score map of head 2
Figure 6: Score map of cross-attentions between input and output patches.
In Fig. 6, we illustrate a score map (12$\times$18) of the cross-attention from the trained CATS. Since both patch length and stride are set to 4, each patch will contain exactly one shock value. We observe that the cross-attentions capture the shocks within the signal and the periodicity of the signal in Fig. 6a and Fig. 6b, respectively. Fig. 6a shows that patches an even number of steps before the current patch contain the shocks of the same direction, resulting in higher attention scores, while odd-numbered steps have lower scores. Moreover, the correlation over 24 steps is clearly demonstrated in patches spaced by multiples of 6 steps, as shown in Fig. 6b. This periodic pattern ensures that the attention mechanism effectively captures the periodicity in ${\mathbf{x}}(t)$, reflecting the model’s ability to leverage this periodic information for more accurate predictions. In Appendix, we provide a detailed explanation.
(a) Forecasting results
(b) Averaged score
(c) L2-H2
(d) L4-H4
Figure 7: Illustration of (a) forecasting result, (b) averaged cross-attention score, and (c,d) patches with the highest score on ETTm1. The score map is averaged from all the heads across layers.
Fig. 7 illustrates (a) forecasting results, (b) a cross-attention score map (5$\times$4) on the ETTm1 dataset, and (c, d) the two pairs with the highest attention scores. We predict 96 steps with input sequence length 96 on ETTm1. The input patches consist of four patches of 24 lengths and one padding patch. As shown in Fig. 7c and 7d, the patches with high attention scores exhibit similar temporal patterns, demonstrating the ability of CATS to detect sequential and periodic patterns.
6 Conclusion
Based on our study, we exploit the advantages of Transformer models in time series forecasting by removing self-attentions and developing a new cross-attention-based architecture. We believe that our model establishes a strong baseline for such forecasting tasks and offers further insights into the complexities of long-term forecasting problems. Our findings provide a reevaluation of self-attentions in this domain, and we hope that future research can critically assess the efficacy and efficiency across various time series analysis tasks. As a limitation, our proposed methods assume channel independence between variables based on the recent work [14]. As the time series data in the real-world are highly correlated, we hope future research can address cross-variate dependency with reduced computation complexity based on the proposed architecture.
7 Acknowledgements
This work was partly supported by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2022-II220984, Development of Artificial Intelligence Technology for Personalized Plug-and-Play Explanation and Verification of Explanation) and the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (No. RS-2024-00338859). This work was also supported by the MSIT(Ministry of Science and ICT), Korea, under the ITRC(Information Technology Research Center) support program (IITP-2024-RS-2024-00438056) supervised by the IITP.
Appendix A Experimental settings
A.1 Datasets
We evaluated the performance using seven datasets commonly used in long-term time series forecasting, including Weather, Traffic, Electricity, ETT (ETTh1, ETTh2, ETTm1, and ETTm2), and M4. These datasets capture a range of periodic characteristics and scenarios that are difficult to predict in the real world, making them highly suitable for tasks, such as long-term time series forecasting, generation, and imputation. Details of these datasets are described in Table 6. The M4 dataset are provided Wu et al. [24], while the remaining datasets are provided in Wu et al. [23].
Table 6: Details of 13 real-world datasets.
Dimension Frequency Timesteps Information Forecasting Horizon
Weather 21 10-min 52,696 Weather (96, 192, 336, 720)
Electricity 321 Hourly 17,544 Electricity (96, 192, 336, 720)
Traffic 862 Hourly 26,304 Transportation (96, 192, 336, 720)
ETTh1 7 Hourly 17,420 Temperature (96, 192, 336, 720)
ETTh2 7 Hourly 17,420 Temperature (96, 192, 336, 720)
ETTm1 7 15-min 69,680 Temperature (96, 192, 336, 720)
ETTm2 7 15-min 69,680 Temperature (96, 192, 336, 720)
M4-Quartely 1 Quartely 48000 Finance 8
M4-Monthly 1 Monthly 96000 Industry 18
M4-Yearly 1 Yearly 46000 Demographic 6
M4-Weekly 1 Weekly 718 Macro 13
M4-Daily 1 Daily 8454 Micro 14
M4-Hourly 1 Hourly 828 Other 48
A.2 Hyperparameter Settings
In every experiment in our paper, following [14], we fixed the random seed of 2021 to enhance the reproducibility of our experiments. Additionally, following numerous studies in the field of time series forecasting [14], we fixed the input sequence length $L=96$. For the forecasting horizon $T$, we also used the widely accepted values, i.e., $[96,192,336,720]$. For our model, in all configurations, we adopt the GeGLU activation function [16] between the two linear layers in the feed-forward network for our model. Additionally, we use learnable positional embedding parameters for the input data and omit positional embeddings for learnable queries to avoid redundant parameter learning.
For the experiments summarized in Table 2 and Table 9, our model uses three cross-attention layers with embedding size $D=256$, number of attention heads $H=32$. Specifically, to avoid overfitting on small datasets [14], we use patch length 48 on the ETTh1 and ETTh2 datasets. Further details on the hyperparameter settings for these experiments are provided in Table 7.
Table 7: Experimental settings used in Table 2 and Table 9.
Metric Layers Embedding Size Query Sharing Input Sequence Length Batch Size Epoch Learning Rate
Weather 3 256 False 96 64 30 $10^{-3}$
Electricity 3 256 False 96 32 30 $10^{-3}$
Traffic 3 256 True 96 32 100 $10^{-3}$
ETTh1 3 256 False 96 256 10 $10^{-3}$
ETTh2 3 256 True 96 256 10 $10^{-3}$
ETTm1 3 256 False 96 128 30 $10^{-3}$
ETTm2 3 256 True 96 128 30 $10^{-3}$
Additionally, for the short-term forecasting experiments on the M4 dataset, we employed a slightly different configuration to better suit the nature of short-term time series. The hyperparameter settings for these experiments are detailed in Table 8. The complete results for these short-term experiments are presented in Table 10, while the full results for the long-term forecasting experiments with a fixed 96 input sequence length are provided in the Appendix in Table 9.
Table 8: Experimental settings of M4 Dataset.
Dataset Layers Embedding Size Input Sequence Length Batch Size Epoch Patience Learning Rate
M4-Quartely 3 64 16 32 30 10 $10^{-3}$
M4-Monthly 3 64 36 32 30 10 $10^{-3}$
M4-Yearly 3 64 12 32 30 10 $10^{-3}$
M4-Weekly 3 64 26 32 30 10 $10^{-3}$
M4-Daily 3 64 28 32 30 10 $10^{-3}$
M4-Hourly 3 128 96 32 30 10 $10^{-3}$
Table 9: Multivariate long-term forecasting results with recent forecasting models and ours for unified hyperparameter settings. The best results are in bold and the second best are underlined.
Models CATS TimeMixer PatchTST Timesnet Crossformer MICN FiLM DLinear Autoformer Informer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Weather 96 0.161 0.207 0.163 0.209 0.186 0.227 0.172 0.220 0.195 0.271 0.198 0.261 0.195 0.236 0.195 0.252 0.266 0.336 0.300 0.384
192 0.208 0.250 0.208 0.250 0.234 0.265 0.219 0.261 0.209 0.277 0.239 0.299 0.239 0.271 0.237 0.295 0.307 0.367 0.598 0.544
336 0.264 0.290 0.251 0.287 0.284 0.301 0.246 0.337 0.273 0.332 0.285 0.336 0.289 0.306 0.282 0.331 0.359 0.395 0.578 0.523
720 0.342 0.341 0.339 0.341 0.356 0.349 0.365 0.359 0.379 0.401 0.351 0.388 0.361 0.351 0.345 0.382 0.419 0.428 1.059 0.741
Electricity 96 0.149 0.237 0.153 0.247 0.190 0.296 0.168 0.272 0.219 0.314 0.180 0.293 0.198 0.274 0.210 0.302 0.201 0.317 0.274 0.368
192 0.163 0.250 0.166 0.256 0.199 0.304 0.184 0.322 0.231 0.322 0.189 0.302 0.198 0.278 0.210 0.305 0.222 0.334 0.296 0.386
336 0.180 0.268 0.185 0.277 0.217 0.319 0.198 0.300 0.246 0.337 0.198 0.312 0.217 0.300 0.223 0.319 0.231 0.443 0.300 0.394
720 0.219 0.302 0.225 0.310 0.258 0.352 0.220 0.320 0.280 0.363 0.217 0.330 0.278 0.356 0.258 0.350 0.254 0.361 0.373 0.439
Traffic 96 0.421 0.270 0.462 0.285 0.526 0.347 0.593 0.321 0.644 0.429 0.577 0.350 0.647 0.384 0.650 0.396 0.613 0.388 0.719 0.391
192 0.436 0.275 0.473 0.296 0.522 0.332 0.617 0.336 0.665 0.431 0.589 0.356 0.600 0.361 0.598 0.370 0.616 0.382 0.696 0.379
336 0.453 0.284 0.498 0.296 0.517 0.334 0.629 0.336 0.674 0.420 0.594 0.358 0.610 0.367 0.605 0.373 0.622 0.337 0.777 0.420
720 0.484 0.303 0.506 0.313 0.552 0.352 0.640 0.350 0.683 0.424 0.613 0.361 0.691 0.425 0.645 0.394 0.660 0.408 0.864 0.472
ETTm1 96 0.318 0.357 0.320 0.357 0.352 0.374 0.338 0.375 0.404 0.426 0.365 0.387 0.353 0.370 0.346 0.374 0.505 0.475 0.672 0.571
192 0.357 0.377 0.361 0.381 0.390 0.393 0.374 0.387 0.450 0.451 0.403 0.408 0.389 0.387 0.382 0.391 0.553 0.496 0.795 0.669
336 0.387 0.401 0.390 0.404 0.421 0.414 0.410 0.411 0.532 0.515 0.436 0.431 0.421 0.408 0.415 0.415 0.621 0.537 1.212 0.871
720 0.448 0.437 0.454 0.441 0.462 0.449 0.478 0.450 0.666 0.589 0.489 0.462 0.481 0.441 0.473 0.451 0.671 0.561 1.166 0.823
ETTm2 96 0.178 0.261 0.175 0.258 0.183 0.270 0.187 0.267 0.287 0.366 0.197 0.296 0.183 0.266 0.193 0.293 0.255 0.339 0.365 0.453
192 0.248 0.308 0.237 0.299 0.255 0.314 0.249 0.309 0.414 0.492 0.284 0.361 0.248 0.305 0.284 0.361 0.281 0.340 0.533 0.563
336 0.304 0.343 0.298 0.340 0.309 0.347 0.321 0.351 0.597 0.542 0.381 0.429 0.309 0.343 0.382 0.429 0.339 0.372 1.363 0.887
720 0.402 0.402 0.391 0.396 0.412 0.404 0.408 0.403 1.730 1.042 0.549 0.522 0.410 0.400 0.558 0.525 0.433 0.432 3.379 1.338
ETTh1 96 0.371 0.395 0.375 0.400 0.460 0.447 0.384 0.402 0.423 0.448 0.426 0.446 0.438 0.433 0.397 0.412 0.449 0.459 0.865 0.713
192 0.426 0.422 0.429 0.421 0.512 0.477 0.436 0.429 0.471 0.474 0.454 0.464 0.493 0.466 0.446 0.441 0.500 0.482 1.008 0.792
336 0.437 0.432 0.484 0.458 0.546 0.496 0.638 0.469 0.570 0.546 0.493 0.487 0.547 0.495 0.489 0.467 0.521 0.496 1.107 0.809
720 0.474 0.461 0.498 0.482 0.544 0.517 0.521 0.500 0.653 0.621 0.526 0.526 0.586 0.538 0.513 0.510 0.514 0.512 1.181 0.865
ETTh2 96 0.287 0.341 0.289 0.341 0.308 0.355 0.340 0.374 0.745 0.584 0.372 0.424 0.322 0.364 0.340 0.394 0.346 0.388 3.755 1.525
192 0.361 0.388 0.372 0.392 0.393 0.405 0.402 0.414 0.877 0.656 0.492 0.492 0.404 0.414 0.482 0.479 0.456 0.453 5.602 1.931
336 0.374 0.403 0.386 0.414 0.427 0.436 0.452 0.452 1.043 0.731 0.607 0.555 0.435 0.445 0.591 0.541 0.482 0.486 4.721 1.835
720 0.412 0.433 0.412 0.434 0.436 0.450 0.462 0.468 1.104 0.763 0.824 0.655 0.447 0.458 0.839 0.661 0.515 0.511 3.647 1.625
Table 10: Full results of univariate short-term forecasting in the M4 dataset. All forecasting horizons are in [6, 48]. The best results are in bold and the second best are underlined.
Models CATS TimeMixer Timesnet PatchTST MICN FiLM DLinear Autoformer Informer
Quarterly SMAPE 9.979 9.996 10.100 10.644 15.214 12.925 12.145 11.338 11.360
MASE 1.164 1.166 1.182 1.278 1.963 1.664 1.520 1.365 1.401
OWA 0.878 0.825 0.890 0.949 1.407 1.193 1.106 1.012 1.027
Monthly SMAPE 12.557 12.605 12.670 13.399 16.943 15.407 13.514 13.958 14.062
MASE 0.916 0.919 0.933 1.031 1.442 1.298 1.037 1.103 1.141
OWA 0.866 0.869 0.878 0.949 1.265 1.144 0.956 1.002 1.024
Yearly SMAPE 13.263 13.206 13.387 16.463 25.022 17.431 16.965 13.974 14.727
MASE 2.967 2.916 2.996 3.967 7.162 4.043 4.283 3.134 3.418
OWA 0.779 0.776 0.786 1.003 1.667 1.042 1.058 0.822 0.881
Others SMAPE 4.560 4.564 4.891 6.558 41.985 7.134 6.709 5.485 24.460
MASE 3.107 3.115 3.302 4.511 62.734 5.09 4.953 3.865 20.960
OWA 0.970 0.982 1.035 1.401 14.313 1.553 1.487 1.187 5.879
Average SMAPE 11.701 11.723 11.829 13.152 19.638 14.863 13.639 12.909 14.086
MASE 1.557 1.559 1.585 1.945 5.947 2.207 2.095 1.771 2.718
OWA 0.838 0.840 0.851 0.998 2.279 1.125 1.051 0.939 1.230
Appendix B Additional Experimental Results
B.1 Performance with Longer Input Sequences
In Table 3, we compared models with the number of parameters, GPU memory consumption, and MSE across different input lengths on ETTm1 with varying input sequence lengths. In this section, we provide comprehensive results on longer input sequence lengths $L=512$. The detailed parameters can be found in Table 11 and the corresponding experimental results are summarized in Table 12. As with unified hyperparameter settings, we follow the settings of the most recent work [21] to ease comparison. Overall, the experimental results clearly illustrate the superiority of CATS over recent forecasting models across multiple datasets and prediction horizons. CATS consistently shows the lowest Mean Squared Error (MSE) and Mean Absolute Error (MAE) across a variety of datasets and forecast horizons. For instance, on Electricity, at the 96 forecast horizon, CATS achieves the best MSE score of 0.144 and the best MAE score of 0.189, underscoring its accuracy in predicting electrical demand.
Table 11: Experimental settings with an input sequence length of 512.
Metric Layers Embedding Size Query Sharing Input Sequence Length Batch Size Epoch Learning Rate
Weather 3 128 False 512 128 30 $10^{-3}$
Electricity 3 128 False 512 32 30 $10^{-3}$
Traffic 3 128 True 512 32 100 $10^{-3}$
ETTh1 3 256 False 512 128 10 $10^{-3}$
ETTh2 3 256 True 512 256 10 $10^{-3}$
ETTm1 3 128 False 512 128 30 $10^{-3}$
ETTm2 3 256 True 512 128 30 $10^{-3}$
Table 12: Multivariate long-term forecasting results with an input sequence length of 512. The best results are in bold and the second best are underlined.
Models CATS TimeMixer PatchTST Timesnet Crossformer MICN FiLM DLinear Autoformer Informer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Weather 96 0.144 0.199 0.147 0.197 0.149 0.198 0.172 0.220 0.232 0.302 0.161 0.229 0.199 0.262 0.176 0.237 0.266 0.336 0.300 0.384
192 0.188 0.240 0.189 0.239 0.194 0.241 0.219 0.261 0.371 0.410 0.220 0.281 0.228 0.288 0.220 0.282 0.307 0.367 0.598 0.544
336 0.238 0.280 0.241 0.280 0.306 0.282 0.246 0.337 0.495 0.515 0.278 0.331 0.267 0.323 0.265 0.319 0.359 0.395 0.578 0.523
720 0.308 0.329 0.310 0.330 0.314 0.334 0.365 0.359 0.526 0.542 0.311 0.356 0.319 0.361 0.323 0.362 0.419 0.428 0.590 0.741
Electricity 96 0.126 0.218 0.129 0.224 0.129 0.222 0.168 0.272 0.150 0.251 0.164 0.269 0.154 0.267 0.140 0.237 0.201 0.317 0.274 0.368
192 0.144 0.235 0.140 0.220 0.147 0.240 0.184 0.322 0.161 0.260 0.177 0.285 0.164 0.258 0.153 0.249 0.222 0.334 0.296 0.386
336 0.159 0.252 0.161 0.255 0.163 0.259 0.198 0.300 0.182 0.281 0.193 0.304 0.188 0.283 0.169 0.267 0.231 0.338 0.300 0.394
720 0.194 0.283 0.194 0.287 0.197 0.290 0.220 0.320 0.251 0.339 0.212 0.321 0.236 0.332 0.203 0.301 0.254 0.361 0.373 0.439
Traffic 96 0.352 0.243 0.360 0.249 0.360 0.249 0.593 0.321 0.514 0.267 0.519 0.309 0.416 0.294 0.410 0.282 0.613 0.388 0.719 0.391
192 0.373 0.253 0.375 0.250 0.379 0.256 0.617 0.336 0.549 0.252 0.537 0.315 0.408 0.288 0.423 0.287 0.616 0.382 0.696 0.379
336 0.387 0.260 0.385 0.270 0.392 0.264 0.629 0.336 0.530 0.300 0.534 0.313 0.425 0.298 0.436 0.296 0.622 0.337 0.777 0.420
720 0.425 0.281 0.430 0.281 0.432 0.286 0.640 0.350 0.573 0.313 0.577 0.325 0.520 0.353 0.466 0.315 0.660 0.408 0.864 0.472
ETTh1 96 0.373 0.401 0.361 0.390 0.370 0.400 0.384 0.402 0.418 0.438 0.421 0.431 0.422 0.432 0.375 0.399 0.449 0.459 0.865 0.713
192 0.401 0.421 0.409 0.414 0.413 0.429 0.436 0.429 0.539 0.517 0.474 0.487 0.462 0.458 0.405 0.416 0.500 0.482 0.080 0.792
336 0.415 0.434 0.430 0.429 0.422 0.440 0.638 0.469 0.709 0.638 0.569 0.551 0.501 0.483 0.439 0.443 0.521 0.496 0.107 0.809
720 0.435 0.446 0.445 0.460 0.447 0.468 0.521 0.500 0.733 0.636 0.770 0.672 0.544 0.526 0.472 0.490 0.514 0.512 0.181 0.865
ETTh2 96 0.256 0.328 0.271 0.330 0.274 0.337 0.340 0.374 0.425 0.463 0.299 0.364 0.323 0.370 0.289 0.353 0.358 0.397 0.755 0.525
192 0.311 0.366 0.317 0.402 0.314 0.382 0.231 0.322 0.473 0.500 0.441 0.454 0.391 0.415 0.383 0.418 0.456 0.453 0.602 0.931
336 0.319 0.382 0.332 0.396 0.329 0.384 0.452 0.452 0.581 0.562 0.654 0.567 0.415 0.440 0.448 0.465 0.482 0.486 0.721 0.835
720 0.395 0.438 0.342 0.408 0.379 0.422 0.462 0.468 0.775 0.665 0.956 0.716 0.441 0.459 0.605 0.551 0.515 0.511 0.647 0.625
ETTm1 96 0.283 0.340 0.291 0.340 0.293 0.346 0.338 0.375 0.361 0.403 0.316 0.362 0.302 0.345 0.299 0.343 0.505 0.475 0.672 0.571
192 0.319 0.363 0.327 0.365 0.333 0.370 0.374 0.387 0.387 0.422 0.363 0.390 0.338 0.368 0.335 0.365 0.553 0.496 0.795 0.669
336 0.351 0.385 0.360 0.381 0.369 0.392 0.410 0.411 0.605 0.572 0.408 0.426 0.373 0.388 0.369 0.386 0.621 0.537 0.212 0.871
720 0.400 0.414 0.415 0.417 0.416 0.420 0.478 0.450 0.703 0.645 0.481 0.476 0.420 0.420 0.425 0.421 0.671 0.561 0.166 0.823
ETTm2 96 0.165 0.256 0.164 0.254 0.166 0.256 0.187 0.267 0.275 0.358 0.179 0.275 0.165 0.256 0.167 0.260 0.255 0.339 0.365 0.453
192 0.221 0.297 0.223 0.295 0.223 0.296 0.249 0.309 0.345 0.400 0.307 0.376 0.222 0.296 0.224 0.303 0.281 0.340 0.533 0.563
336 0.274 0.334 0.279 0.330 0.274 0.329 0.321 0.351 0.657 0.528 0.325 0.388 0.277 0.333 0.281 0.342 0.339 0.372 0.363 0.887
720 0.362 0.390 0.359 0.383 0.362 0.385 0.408 0.403 0.208 0.753 0.502 0.490 0.371 0.389 0.397 0.421 0.422 0.419 0.379 0.388
B.2 Additional Results for Section 5.2
Table 13: Comparison of models with the number of parameters, GPU memory consumption, and MSE across different input sequence lengths on ETTm1. Full results of Table 3.
Parameters across different input lengths
Models 96 192 336 512 720 1440 2880
PatchTST 1,506,384 2,612,304 4,271,184 6,298,704 8,694,864 16,989,264 33,578,064
TimeMixer 190,313 484,217 1,129,193 2,250,137 4,046,633 14,211,593 52,912,313
DLinear 139,680 277,920 485,280 738,720 1,038,240 2,075,040 4,148,640
CATS 360,264 360,776 361,544 362,440 363,592 367,432 375,112
GPU Memory Consumption across different input lengths
Models 96 192 336 512 720 1440 2880
PatchTST 2,234MB 2,650MB 3,484MB 4,914MB 7,368MB 2,1968MB 58,590MB
TimeMixer 2,204MB 2,522MB 2,914MB 3,414MB 3, [^8] MB 5,876MB 10,324MB
DLinear 1,098MB 1,102MB 1,104MB 1,114MB 1,114MB 1,154MB 1,214MB
CATS 1,712MB 1,796MB 1,884MB 2,042MB 2,140MB 2,700MB 3,826MB
MSE across different input lengths
Models 96 192 336 512 720 1440 2880
PatchTST 0.457 0.424 0.418 0.420 0.418 0.420 0.413
TimeMixer 0.454 0.433 0.428 0.436 0.425 0.414 0.472
DLinear 0.473 0.438 0.426 0.427 0.422 0.401 0.408
CATS 0.450 0.418 0.407 0.400 0.402 0.399 0.395
Table 14: Comparison of models with the number of parameters, GPU memory consumption, and MSE across different input sequence lengths on Weather.
Parameters across different input lengths
Models 96 192 336 512 720 1440 2880
PatchTST 1,506,384 2,612,304 4,271,184 6,298,704 8,694,864 16,989,264 33,578,064
TimeMixer 219,249 595,305 1,465,569 3,028,185 5,582,529 20,343,969 77,423,049
DLinear 139,680 277,920 485,280 738,720 1,038,240 2,075,040 4,148,640
CATS 370,344 370,856 371,624 372,520 373,672 377,512 385,192
GPU Memory Consumption across different input lengths
Models 96 192 336 512 720 1440 2880
PatchTST 1,680MB 2,110MB 2,762MB 4,596MB 5,726MB 16,472MB 45,278MB
TimeMixer 1,894MB 2,154MB 2,728MB 3,414MB 4,356MB 8,358MB 20,624MB
DLinear 1,106MB 1,114MB 1,188MB 1,188MB 1,188MB 1,362MB 1,632MB
CATS 1,522MB 1,590MB 1,665MB 1,755MB 1,892MB 2,282MB 3,140MB
MSE across different input lengths
Models 96 192 336 512 720 1440 2880
PatchTST 0.351 0.336 0.320 0.315 0.309 0.308 0.312
TimeMixer 0.339 0.331 0.318 0.319 0.324 0.318 0.327
DLinear 0.346 0.334 0.325 0.320 0.316 0.311 0.309
CATS 0.342 0.325 0.314 0.308 0.305 0.301 0.291
We provide additional experimental results to support the findings discussed in Section 5.2. Tables 13 and 14 summarize detailed comparisons of the number of parameters, GPU memory consumption, and MSE across different input lengths for the ETTm1 and Weather datasets, respectively. The linear models, TimeMixer and DLinear, exhibit smaller parameters for shorter input lengths. Despite CATS having slightly more parameters than TimeMixer for smaller inputs, it outperforms in terms of memory usage and MSE. This suggests that CATS is more efficient and effective in handling shorter inputs. For PatchTST, the number of parameters does not increase within the actual Transformer backbone as the input length increases. However, due to the need to flatten and project all inputs at the end, the parameters scale linearly with the input length. This highlights a limitation of the Encoder’s architecture. On the other hand, TimeMixer’s parameters grow almost quadratically as the input length doubles. Similarly, DLinear’s parameters increase linearly with the input length. Our proposed model, CATS demonstrates significant efficiency through parameter sharing, where the parameters hardly increase with longer inputs. Notably, from an input length of 336, CATS has fewer parameters than DLinear, showcasing the deep learning model’s advantage in detecting inherent patterns in the data.
Regarding GPU memory consumption, we observe that both PatchTST and TimeMixer require significantly more GPU memory as the input length increases. For example, PatchTST’s GPU memory usage scales drastically, making it less feasible for long input sequences. TimeMixer also shows an increase in GPU memory consumption, although it is less severe than PatchTST. In contrast, DLinear maintains a relatively constant GPU memory usage, demonstrating its efficiency in terms of computational resources. However, CATS stands out by offering a balanced approach, with moderate GPU memory usage that scales more favorably compared to PatchTST and TimeMixer. This balance between memory efficiency and performance is crucial for practical applications requiring long-term time series forecasting.
Furthermore, when analyzing the MSE across different input lengths, CATS consistently shows the best performance. It maintains lower MSE compared to other models across all input lengths. This robustness in performance, combined with its efficient parameter and memory usage, highlights the superiority of CATS in long-term time series forecasting tasks. Overall, these results show the advantages of CATS in terms of parameter efficiency, GPU memory consumption, and forecasting accuracy. These findings support the proposed model’s potential for practical and scalable time series forecasting solutions.
Table 15 presents the full results on the Traffic dataset. Here, we use the Traffic dataset with a batch size of 8. All GPU memory consumption was measured in a setting using four multi-GPUs. As shown in Table 15, CATS with a 2880 input sequence length consistently outperforms models with a 512 input sequence length, including PatchTST and TimeMixer. Specifically, CATS demonstrates fewer parameters, lower GPU memory consumption, and faster running speeds. These results highlight the efficiency of CATS with large input sizes. The Traffic dataset, characterized by high-dimensional data, shows a significant reduction in MSE when using longer input sequences.
Table 16 provides the full results on the Electricity dataset. Similar to the Traffic dataset, CATS shows superior efficiency in training, particularly with an input size of 2880, across all cases. Here, we use the Electricity dataset with a batch size of 32. All GPU memory consumption was measured in a setting using four multi-GPUs. In this experiment, CATS with a 512 input sequence length did not use parameter sharing for queries, while CATS with a 2880 input sequence length did. This demonstrates the effectiveness of query parameter sharing when utilizing large amounts of data for training. The results confirm that query sharing among dimensions leads to greater efficiency and improved performance.
Table 15: Comparison of models with the number of parameters, GPU memory consumption, running speed, and MSE across different forecasting horizon sizes on Traffic. Full results of Fig. 5.
Horizon Models Paramters Gpu Memory Running Time MSE
96 PatchTST 1,186,272 28.54GB 0.1390s/iter 0.360
TimeMixer 2,442,961 38.12GB 0.2548s/iter 0.360
CATS ($L=512$) 357,496 5.81GB 0.0533s/iter 0.352
CATS ($L=2880$) 370,168 9.79GB 0.1158s/iter 0.339
192 PatchTST 1,972,800 28.34GB 0.1412s/iter 0.379
TimeMixer 2,535,505 38.13GB 0.2596s/iter 0.375
CATS ($L=512$) 357,592 6.73GB 0.0571s/iter 0.373
CATS ($L=2880$) 370,264 11.10GB 0.1209s/iter 0.362
336 PatchTST 3,152,592 28.91GB 0.1487s/iter 0.392
TimeMixer 2,674,321 38.69GB 0.2647s/iter 0.385
CATS ($L=512$) 357,736 7.46GB 0.0584s/iter 0.387
CATS ($L=2880$) 370,408 12.72GB 0.1266s/iter 0.379
720 PatchTST 6,298,704 29.15GB 0.1628s/iter 0.432
TimeMixer 3,044,497 41.17GB 0.2777s/iter 0.430
CATS ($L=512$) 358,120 10.10GB 0.0734s/iter 0.423
CATS ($L=2880$) 370,792 18.40GB 0.1556s/iter 0.420
Table 16: Comparison of models with the number of parameters, GPU memory consumption, running speed, and MSE across different forecasting horizon sizes on Electricity.
Horizon Models Paramters Gpu Memory Running Time MSE
96 PatchTST 1,186,272 40.36GB 0.2021s/iter 0.129
TimeMixer 2,429,049 33.80GB 0.2118s/iter 0.129
CATS ($L=512$) 388,216 6.89GB 0.0587s/iter 0.126
CATS ($L=2880$) 370,168 12.82GB 0.1653s/iter 0.126
192 PatchTST 1,972,800 40.39GB 0.2048s/iter 0.147
TimeMixer 2,521,593 33.81GB 0.2212s/iter 0.140
CATS ($L=512$) 419,032 8.07GB 0.0636s/iter 0.144
CATS ($L=2880$) 370,264 14.70GB 0.1725s/iter 0.139
336 PatchTST 3,152,592 40.42GB 0.2070s/iter 0.163
TimeMixer 2,660,409 34.24GB 0.2314s/iter 0.161
CATS ($L=512$) 465,256 9.15GB 0.0690s/iter 0.159
CATS ($L=2880$) 370,408 17.38GB 0.1839s/iter 0.153
720 PatchTST 6,298,704 41.40GB 0.2313s/iter 0.197
TimeMixer 3,030,585 36.13GB 0.2478s/iter 0.194
CATS ($L=512$) 588,520 12.77GB 0.0964s/iter 0.194
CATS ($L=2880$) 370,792 25.86GB 0.2262s/iter 0.183
B.3 Ablation Study on Query-adaptive Masking
In this section, we demonstrate the effectiveness of query-adaptive masking compared to dropout, which is a widely adopted technique in transformer-based forecasting models. We consider four different setups: using only dropout, using query-adaptive masking with fixed probabilities, query-adaptive masking with linearly increasing probabilities, and using both methods simultaneously. As shown in Fig. 8, the query-adaptive masking shows better forecasting performance and faster converge speed compared to dropout. Applying a gradually increasing masking probability based on the horizon predicted by the query shows slight performance improvements over using a fixed probability or combining with dropout. In contrast, using dropout alone shows noticeable differences in both convergence speed and overall performance. This demonstrates that when multiple inputs with different forecasting horizons share a single model, probabilistic masking is more beneficial for model training than dropout.
(a) ETTm1 with $T=720$
(b) Weather with $T=720$
Figure 8: Comparison of performance with query-adaptive masking with two different probabilities, dropout, and using both query-adaptive masking and dropout. The results of $p=0.1$ to $0.7$ indicate a probability masking that is linearly increased proportionally to the horizon predicted by the query.
Appendix C Detailed Explanation of Section 5.5
In this section, we provide the detailed results of experiments in Section 5.5. We first restate the formulation of two independent signals used in Section 5.5 as follows:
${\mathbf{x}}(t)$ =$=\{x_{(t\bmod\tau)}\}^{\infty}_{t=1},\quad x_{i}\sim\mathcal{N}(0,1)\quad(i=0,1,\ldots,\tau-1),$
${\bm{y}}(t)$ =$=\begin{cases}+k&\text{if }t\equiv 0\pmod{S}\\ -k&\text{if }t\equiv\frac{1}{2}S\pmod{S}\end{cases},$
We use the model parameters as follows: the patch length is 4 without overlapping, the decoder has 1 layer, and there are 2 attention heads. The signals ${\mathbf{x}}(t)$ and ${\bm{y}}(t)$ are defined with $\tau=24$, $S=8$, and $k=5$. The visualization of synthetic data is shown in Fig. 9. We utilize an input sequence length $L=48$ and a forecasting horizon $T=72$. This setup allows us to generate time series data with distinct periodic components.
(a) Two input signals
(b) Sum of two input signals
Figure 9: Visualization of input signals for toy experiment.
In the main paper, Fig. 6 displays a cross-attention score map between the input patch and the output patch derived from this experiment. The left figure presents the attention score of the first attention head, illustrating the model’s ability to detect shocks within the signal. The right figure more clearly demonstrates the periodicity of the signal. Given that the patch length and stride are both set to 4, each patch will contain exactly one shock value, either -5 or +5. This is because the shocks occur every 4 steps, alternating between positive and negative shocks. Consequently, the patch immediately preceding the current patch will contain a different shock, leading to lower attention scores due to the differing shock values. In contrast, patches that are an even number of steps before the current patch will contain the same type of shock, resulting in higher attention scores. These points are well illustrated in Fig. 6a, where the varying attention scores correspond to the presence of alternating shocks. This pattern helps to highlight the alternating shock signal within the data.
Additionally, if there is a correlation with the series preceding 24 steps, the patches that are 6 steps or multiples of 6 steps before the current patch will exhibit high attention scores due to the periodic nature of the signal ${\mathbf{x}}(t)$. The diagonal formation of the attention scores, which accurately follows a period of 24, is clearly depicted in Fig. 6b, highlighting the model’s capability to utilize fixed-period input patches to predict future outcomes. This periodic pattern ensures that the attention mechanism effectively captures the 24-step periodicity in ${\mathbf{x}}(t)$, reflecting the model’s ability to leverage this periodic information for more accurate predictions.
This experimental configuration provides a robust framework to evaluate how well our proposed model captures and interprets the underlying patterns in the data, specifically focusing on the alternating shock signal and the periodic nature of the normal signal. This dual emphasis on both the shock signal and the periodicity of the normal signal enhances the interpretability and predictive performance of the model, distinctly demonstrating how the model leverages periodic information to enhance prediction accuracy.
To push further, we reproduce the experiment of Fig. 7 with other datasets used in forecasting tasks. We illustrate the results of the Weather, Traffic, Electricity, ETTm2, ETTh1, and ETTh2 in Figures 10, 11, 12, 13, 14, and 15, respectively. For each figure, (a) represents the forecasting results, (b) shows the cross-attention score map, and (c) and (d) illustrate the two pairs with the highest attention scores. For all figures, our attention-based explanation successfully discovers similar periodic patterns. Therefore, we believe that our model has the potential to provide a clearer understanding of the mechanisms underlying forecasting predictions. We hope that future research will continue to explore and expand upon this foundation.
(a) Forecasting results
(b) Averaged score
(c) L1-H2
(d) L1-H3
Figure 10: Illustration of (a) forecasting result, (b) averaged cross-attention score, and (c,d) patches with the highest score on Weather. The score map is averaged from all the heads across layers.
(a) Forecasting results
(b) Averaged score
(c) L4-H3
(d) L1-H4
Figure 11: Illustration of (a) forecasting result, (b) averaged cross-attention score, and (c,d) patches with the highest score on Traffic. The score map is averaged from all the heads across layers.
(a) Forecasting results
(b) Averaged score
(c) L1-H2
(d) L1-H3
Figure 12: Illustration of (a) forecasting result, (b) averaged cross-attention score, and (c,d) patches with the highest score on Electricity. The score map is averaged from all the heads across layers.
(a) Forecasting results
(b) Averaged score
(c) L1-H1
(d) L2-H2
Figure 13: Illustration of (a) forecasting result, (b) averaged cross-attention score, and (c,d) patches with the highest score on ETTm2. The score map is averaged from all the heads across layers.
(a) Forecasting results
(b) Averaged score
(c) L2-H1
(d) L2-H2
Figure 14: Illustration of (a) forecasting result, (b) averaged cross-attention score, and (c,d) patches with the highest score on ETTh1. The score map is averaged from all the heads across layers.
(a) Forecasting results
(b) Averaged score
(c) L1-H1
(d) L1-H2
Figure 15: Illustration of (a) forecasting result, (b) averaged cross-attention score, and (c,d) patches with the highest score on ETTh2. The score map is averaged from all the heads across layers.